# Complex magnetic landscape in kagome metal TbTi3Bi4 and competition with family of Ln2-xTi6+xBi9 (Ln: Tb–Lu) shurikagome metals

### v1

@@ III.1 Crystal Structure and Synthesis
As alluded to before, TbTi3Bi4 is not stable under conditions presented in our first manuscript.[28] Samples of large rare-earth ATi3Bi4 compounds are readily grown through a simple flux (cooled from 1100°C to 600°C at a rate of 2°C/hr). However, samples containing Tb will exclusively form Tb2-xTi6+xBi9. Figure 2(a) shows a small section of the Tb–Ti–Bi phase diagram near several of the phases of interest. Data for the existing binary and pseudobinary lines, including the single phase region surrounding Ti8Bi9 were adapted from published Ti–Bi, Tb–Bi binary diagrams.[41, 42] Note the proximity of TbTi3Bi4 and Tb2-xTi6+xBi9 (particularly the proximity of the the hypothetically ideal Tb2Ti6Bi9). The large single-phase region of Ti8Bi9 (and the associated alloys with Tb2-xTi6+xBi9) is problematic, as it isolates a large swath of the phase diagram from Bi-rich fluxes. The single phase region even persists beyond where a hypothetical Bi2Ti6Bi9 exists (complete substitution of Bi on the mixed sites). There have also been reports that oxygen atoms can substitute in the center of the Ti4 tetrahedra, creating compositions like Ti8Bi9O0.25.[43] However, if Tb2Ti6Bi9 is not stable, then TbTi3Sb4 should still be accessible from a Bi self-flux.
We designed a series of experiments to test both the thermal stability of Tb2-xTi6+xBi9 and compositional dependence. We first observed that samples allowed to completely freeze (T=270°C) often exhibited Tb2-xTi6+xBi9 crystals that were rounded, as though they started to decompose. Modifying the flux towards Tb-rich conditions instead resulted in rounded crystals of TbBi. In both cases, the corners of the decomposing crystals were encased in agglomerates of TbTi3Bi4. This highlighted that TbTi3Bi4 is a low-temperature phase. Subsequent quenching experiments further suggested that TbTi3Bi4 is likely only stable below 600°C. Samples grown above 600°C reveal two alkemade triangles; 1) three phase TbBi–Tb2-xTi6+xBi9–Bi and 2) two phase Tb2-xTi6+xBi9–Bi. Samples grown below 600°C reveal instead; 1) three phase TbBi–TbTi3Bi4–Bi and 2) three phase TbTi3Bi4–Tb2-xTi6+xBi9–Bi. Figure 2(b) presents a simple projection of Figure 2(a) along the TiBi–TbBi pseudobinary line, summarizing the relative proximity and temperature stability of these results.
@@
Crystals of TbTi3Sb4 are 1-5 mm pseudo-hexagonal crystals with a brilliant metallic luster. They are readily exfoliated and are lightly air-sensitive. Samples can be worked with in air, though they will swell and spall under extended (1 day) exposure. Crystals of Ln2-xTi6+xBi9 are typically truncated square plates (nearly octagonal) with a dark gray luster. Samples are readily cleaved with a blade parallel to the ab-plane. Ln2-xTi6+xBi9 crystals are substantially more sensitive to air, and crystals will completely disintegrate within a few hours. In both cases, the samples expand and self-cleave along the ab-plane upon extended exposure to air.
@@ III.2 Properties of TbTi3Bi4
Due to the reduced symmetry (Fmmm), the ATi3Bi4 materials tend to be highly anisotropic, particularly with regards to the quasi-1D zig-zag chains. TbTi3Bi4 is no exception. To diagnose the potential anisotropies in the material, we first begin with a study of the magnetization along different principle crystallograpic axes. Figure 3(a) presents the magnetization of a single crystal of TbTi3Bi4 with the magnetic field parallel to the [100], [010], and [001] directions. TbTi3Bi4 appears to adopt an antiferromagnetic ground state, with two primary transitions at 24 K and 3 K. The [100] direction corresponds to the “easy axis,” and orienting the magnetic field parallel to the [010] or [001] induces a substantially weaker effect. The behavior is reminiscent of GdTi3Bi4[28], though crystals of TbTi3Bi4 respond negligibly to fields oriented along [001].
@@
Figure 3(b) demonstrates the isothermal magnetization of TbTi3Bi4 with the magnetic field parallel to the [100], [010], and [001] directions. Upon cursory inspection, the data again resembled that of GdTi3Bi4 (two metamagnetic transitions when ). However, the metamagnetic transitions in TbTi3Bi4 exhibit substantial field hysteresis. The data saturates at approximately The first transition plateaus at approximately 9.2 , in good agreement with the 9 expected for Tb3+. The first metamagnetic transition first plateaus at a moment of 2.9-3.2 , which is approximately 1/3 of the expected net moment. Furthermore, closer inspection at the second metamagnetic transition reveals an unusual inflection midway through the rising field direction (see Figure 3(b,inset)). This inflection occurs at approximately 6.2 slightly above 2/3 the expected moment. This inflection and other properties of the isothermal magnetization curves will be studied in detail below. Similar to the temperature-dependent magnetization, there are only weak, linear responses when or . Figure 3(c) provides a series of magnetization curves collected while rotating between , , and in 15°increments. The angular dependence is relatively simple, with the transitions being stretched towards higher fields with increasing deviations from , again suggesting strong easy-plane (and further easy-[100]) anisotropy.
@@
Figure 3(c) presents the inverse susceptibility, collected under a 100 Oe field where . The high temperature data exhibits a slight upwards curvature, potentially due to a temperature-independent contribution from core diamagnetism (bismuth) or the quartz sample holder. We show Curie-Weiss fits to the data with (gray) and without (black) a temperature-independent term. While the gray curve fits over a longer range of data, it requires a large emu mol-1 Oe-1 and yields an unphysical effective moment of 11 . We opt instead to utilize a limited high-temperature regime (red) for a simpler Curie-Weiss fit, which yields a =53 K and a = 9.9. The effective moment agrees well with that expected of Tb3+ (9.72), though the Weiss temperature is unphysically large (53 K), likely affected by the temperature independent contribution, limited temperature range, and crystal-field effects.
@@
As a consistency check with the magnetization data, we performed zero-field heat capacity measurements down to 100 mK. The sample was initially oriented with such that measurements could also be made for the intermediate plateaus seen in Figure 2(b). The first order hysteretic nature of the transitions seen in the isothermal magnetization made the quantative analysis of field-dependent heat capacity challenging, and we have opted to reserve this data for a later study. Instead, we focus on the analysis of the zero-field heat capacity utilizing samples of LaTi3Bi4 as a nonmagnetic standard. Figure 3(d) demonstrates the experimental heat capacity data (green) as a function of temperature, plotted in . The two magnetic transitions are evident around 20 K and 2 K. At temperatures below 2 K, a nuclear Schottky peak dominates the heat capacity signal. The peak is still increasing at 0.2 K, which agrees relatively well with a Tb-based Schottky peak ( K).[44] Utilizing LaTi3Bi4 as the nonmagnetic standard (black) and invoking a simple nuclear Schottky model , we can extract the magnetic entropy for TbTi3Bi4.
@@
Figure 3(e) shows the integrated magnetic entropy (green) as a function of temperature. The extracted magnetic contribution to the heat capacity () is also shown (gray) as a schematic reference for the magnetic transition temperatures (not to scale). Interestingly, integrating up to the primary magnetic transition appears to correspond with an entropy release of approximately . Contributions to the magnetic entropy gradually continue up to 100 K, however, consistent with observations in other ATi3Bi4 materials.[28] By 100 K we recover approximately 90% of the total entropy expected for Tb3+. Considering that the higher temerature heat capacity data tends to be noisier (and that the nonmagnetic subtraction is more prone to error at high temperatures), this is remarkably good agreement with expectations.
We now return to the hysteretic metamagnetic transitions observed in Figure 3(b). Closer inspection into the second metamagnetic transition reveals a complex magnetic landscape which can only be elucidated with high density phase maps. Figure 4(a,d) present our efforts to map the magnetic landscape of TbTi3Bi4, demonstrating a set of isothermal magnetization curves collected in both the increasing (a) and decreasing (d) field directions. One can immediately see that the sharpness of the second transition belies a network of subtle changes in the magnetization. Taking the derivative of the magnetization data shown in (a,d) yields the phase diagrams shown in Figure 4(b,e). Immediately one can see that the plateau visible in the 1.8 K magnetization data shown before in Figure 3(b) fragments into a multitude of unique states at intermediate temperatures from 2–15 K. To better highlight the intricacy of the magnetic landscape near the potential 2/3 magnetization plateau, we collected higher density data over the limited ranges highlighted by the black dashed lines.
@@
Figure 4(c,f) demonstrate the phase diagram for TbTi3Bi4 in the vicinity of the second metamagnetic transition. At least 5 unique phase pockets can be observed. Furthermore, while the number of pockets appears to be roughly conserved, the qualitative behavior of the diagram depends strongly on whether the data is collected upon increasing or decreasing field. Data was also collected for the zero-field-cooled virgin curves, though they are qualitatively similar to Figure 4(a,b,c) and have been omitted for brevity. A series of line cuts (Figure 4(c,right)) through the phase diagram data (white dashed lines) helps to clarify the clear boundaries between the phase pockets. Considering the complexity of the magnetic landscape, neutron diffraction experiments will be invaluable to identify the types and natures of the individual magnetic phases.
Antiferromagnetic metals are not particularly common, particularly so for materials that can be readily exfoliated. While the low temperature magnetism of the rare-earth metals prevents facile applications within fields like spintronics, understanding how the magnetism interacts with the transport properties is key to understanding the connection between the kagome network (which dominates the electronic states near E) and the magnetism. Figure 5(a) shows a basic resistivity curve as a function of temperature under zero-field conditions. Our experiment passes current parallel to the [010] direction, and orients the field such that (oriented to induce the metamagnetic transitions). The crystal quality is good (RRR40) and the zero-field transport is indicative of a relatively high mobility metal. The magnetic transition at 20 K is evident in the resistivity, highlighted by a sharp cusp and subsequent drop. This is the expected result for ordering from the paramagnetic regime, sharply reducing spin-scattering.
@@
We now apply a magnetic field to examine the magnetoresistance (MR) of TbTi3Bi4 in the same experimental setup. Figure 5(b) shows the symmetrized transverse magnetoresistance with and at 1.8 K. As one would expect from the magnetization results, we observe two sharp transitions in the magnetoresistance, one for each of the metamagnetic features. The magnetoresistance retains the qualitative field hysteresis of the magnetization data, both in magnitude and location. The first metamagnetic transition marks a sharp increase in the magnetoresistance of approximately 30–40%. The second metamagnetic transition marks a sharp decrease in the magnetoresistance to approximately 50%, which one may expect for a total polarization of the Tb moments. We find the sharp switching between positive and negative MR intriguing in TbTi3Bi4, particularly for a system which is already quite metallic (1Ohm-cm at 1.8 K).
@@
Given the strong MR response in TbTi3Bi4, we were curious to investigate whether the complex landscape of magnetic transitions seen in the phase diagrams would influence the electronic transport. Figure 5(c) provides a series of magnetoresistance curves as a function of temperature from 2 K to 21 K in 1 K steps. For simplicity and clarity, we are only showing the data under increasing fields. We would like to highlight the behavior near the second transition of the symmetrized data (3 T). Immediately before the transition to negative magnetoresistance, we observe that the magnetoresistance cusp becomes a peak – then splits into two peaks – merge back together – and then vanish as the system approaches the Neel temperature. These effects occur at the intermediate temperatures consistent with the appearance of the various phase pockets in Figure 4(c). This further suggests that these are bulk effects corresponding to distinct types of magnetic order.
The sample geometry prevents us from measuring the Hall response where and . The in-plane moments would require thinning the sample perpendicular to the natural layers of the crystal, which typically causes the sample to flake and fall apart. However, we can readily orient the sample such that and to explore the Hall effect and magnetoresistance when the metamagnetic transitions are not induced. Figure 5(d) shows the antisymmetrized Hall resistivity for TbTi3Bi4 over two temperature regimes. The top figure shows the Hall response from 2 K to 18 K, and the bottom figure highlights temperatures above the magnetic transition temperature (20 K to 30 K). No subtractions or corrections have been performed besides antisymmetrization. All data sets exhibit a strongly nonlinear Hall response over all fields. A signal that looks deceptively like an anomalous Hall signal is evident at 2 K, though the signal persists in shape and magnitude above the magnetic transition temperature. Furthermore, at high temperatures we recover an additional inflection in the data. The signal is also relatively small, leading us to suspect that we may be seeing the effect of multiband transport. We can look to some prior work on NdTi3Bi4 for some guidance – since the electronic structure of these materials doesn’t appreciably change with the choice of rare-earth. Experimental ARPES data has observed a Dirac-like feature near the Fermi level[38, 37, 35, 36] and has further observed gapping of the Dirac-like bands as NdTi3Bi4 transitions into the magnetic ground state[35]. The sample exhibits strong positive magnetoresistance (200%) when and , in stark contrast to behavior when and .

### v2

@@ III.1 Crystal Structure and Synthesis
As alluded to before, TbTi3Bi4 does not appear to form under conditions presented in our first manuscript.[31] Samples of large rare-earth LnTi3Bi4 compounds are readily grown through a bismuth self-flux cooled from 1100°C to 600°C. However, reactions utilizing Tb will exclusively form Tb2-xTi6+xBi9. To help understand the shift towards crystallization of the Tb2-xTi6+xBi9 phase, we first turned to examine the known phases in the vicinity of TbTi3Bi4. Figure 2(a) shows a small section of the Tb–Ti–Bi phase diagram near several of the phases of interest. Data for the existing binary and pseudobinary lines, including the single phase region surrounding Ti8Bi9 were adapted from published Ti–Bi, Tb–Bi binary diagrams.[51, 52] The large Ti–Bi offstoichiometry in samples of Ti8Bi9 is remarkable, and suggests that solid solubility exits beyond where the hypothetical compound Bi2Ti6Bi9 exists (written in this form to highlight substitution of Bi on the mixed sites). There have also been reports that oxygen atoms can substitute in the center of the Ti4 tetrahedra, creating compositions like Ti8Bi9O0.25.[53] Such observations are consistent with the structural flexibility needed to host the Ln disorder proposed here.
It appears that the fully-occupied Tb2Ti6Bi9 is not stable under conditions explored here, as flux growths containing different concentrations of Ln:Ti:Bi consistently produce samples of Tb2-xTi6+xBi9 which are sub-occupied. This is fortunate from the perspective of TbTi3Bi4, as the presence of Tb2Ti6Bi9 would hinder our ability to form TbTi3Bi4 from a bismuth-rich (low-temperature melting point) mixture. Given the absence of Tb2Ti6Bi9, we hypothesized that TbTi3Bi4 could potentially be accessible from a Bi self-flux. However, if the phase exists, it must be a low-temperature phase, which is out-competed by the thermal stability of Tb2-xTi6+xBi9.
One preliminary observation that supported our hypothesis was the rounding (resorption) of the corners and edges of Tb2-xTi6+xBi9 crystals at temperatures near 600°C (flux composition 2:3:20). Samples quenched at higher temperatures (e.g. 800°C) showed nicely faceted Tb2-xTi6+xBi9 samples with no resorption, suggesting that the shurikagome is either: 1) unstable at low temperatures, or 2) is consumed by crystallization of a low-temperature phase. A similar phenomenon was noted with more Tb-rich compositions (flux composition 5:5:20) and the high melting point TbBi binary. Given sufficient time at low-temperatures, the decomposed corners of Tb2-xTi6+xBi9 became encased in recrystallized agglomerates of TbTi3Bi4. This highlighted that TbTi3Bi4 is a low-temperature phase and evolves from the dissolution of the high-temperature compounds. Subsequent quenching experiments further suggested that TbTi3Bi4 is likely only stable below 600°C.
Thus, we propose that TbTi3Bi4 is a low-temperature phase with a maximum stability of approximately 600°C in a bismuth self-flux. Upon cooling, it is preceded by either TbBi or Tb2-xTi6+xBi9, depending on the flux composition. To circumvent formation of the competing phases, we quench samples from 1100° C to 600°C. This step prevents extended nucleation and growth of Tb2-xTi6+xBi9. The samples are then slow-cooled from 600°C to 400°C at a rate of 1°C/hr, ensuring resorption of any shurikagome nuclei and encouraging growth of TbTi3Bi4. Batches grown in this manner are primarily composed of nicely-faceted TbTi3Bi4. We are pleased to see that these observations agree nicely with recent results from other groups.[54, 55]
@@
Crystals of TbTi3Bi4 are 1–5 mm pseudo-hexagonal crystals with a brilliant metallic luster. They are extremely soft, readily exfoliated parallel to the ab plane, and are lightly air-sensitive. In the antimonide analog, we have observed that both M-Sb and Ln-Sb terminations can be achieved through exfoliation, and we expect similar results for the bismuthide.[32] Crystals of Ln2-xTi6+xBi9 are typically truncated square plates (nearly octagonal) with a dark gray luster. Crystals are substantially more brittle than the LnTi3Bi4 compounds but can be readily cleaved parallel to the ab-plane (perpendicular to the Ln chains) with a sharp blade. Ln2-xTi6+xBi9 crystals are substantially more sensitive to air, and crystals will completely disintegrate within a few hours. Curiously, both samples decompose in the same manner. Upon extended exposure, the samples swell substantially and spall layer-by-layer.
@@ III.2 Properties of TbTi3Bi4
Due to the reduced symmetry (Fmmm), the LnTi3Bi4 materials tend to be highly anisotropic, particularly with regards to the quasi-1D zig-zag chains. TbTi3Bi4 is no exception. To diagnose the potential anisotropies in the material, we first begin with a study of the magnetization along different principle crystallograpic axes. Figure 3(a) presents the field-cooled (FC) magnetization curve of a single crystal of TbTi3Bi4 with the magnetic field (100 Oe) parallel to the [100], [010], and [001] directions. At these low fields, there is no difference between the field-cooled (FC) and zero-field-cooled (ZFC) measurements. TbTi3Bi4 appears to adopt an antiferromagnetic ground state, with two transitions at 24 K and 3 K. The [100] direction corresponds to the “easy axis,” and orienting the magnetic field parallel to the [010] or [001] induces a substantially weaker effect.
Figure 3(b) demonstrates the isothermal magnetization of TbTi3Bi4 with the magnetic field parallel to the [100], [010], and [001] directions at 1.8 K. Due to the hysteretic nature of TbTi3Bi4, care must be taken to control the magnetic and thermal history of the samples. For example, Figure 3(b) () shows three different field paths: 1) ZFC virgin magnetization data (gray), 2) decreasing field (purple, down arrow), and 3) increasing field (purple, up arrow). The virgin curve is only obtained upon the first field sweep after cooling below 24 K with no applied field (ZFC). All subsequent field sweeps will fall onto the purple data, tracing either the field increasing (up arrow) or field decreasing (down arrow) path. Continuing the data into the negative field direction is not necessary as the system exhibits no coercivity at zero-field, and loops collected from 0 T to 7 T are functionally identical to those from 0 T to 7 T. For the vast majority of our analysis we will omit the virgin magnetization curves. The differences between the ZFC virgin curve and the next positive field sweep are likely due to the domain structure that forms on cooling in zero field, and are not the focus of this work.
@@
The isothermal magnetization data saturates at approximately 9.2 by 3 T, in good agreement with the M expected for Tb3+. The first metamagnetic transition plateaus at a moment of 2.9 –3.2 , which is approximately 1/3 M. Closer inspection midway through the second metamagnetic transition reveals an unusual inflection (see Figure 3(b,inset)). This inflection occurs at approximately 6.2 , slightly above 2/3 M. This inflection and other properties of the isothermal magnetization curves will be studied in the following figure. As implied by the temperature-dependent magnetization shown in Figure 3(a), the magnetic response is suppressed quickly as we rotate the field away from the [100] direction. Figure 3(c) provides a series of magnetization curves collected while rotating between , , and in 15° increments. Only the increasing and decreasing field directions are shown (no ZFC virgin magnetization curve). The angular dependence is relatively simple, with the transitions being stretched towards higher fields with increasing deviations from , again suggesting strong easy-plane (and further, easy-[100]) anisotropy.
@@
Figure 3(d) presents the inverse susceptibility, collected under a 100 Oe (FC) field where . The high temperature data exhibits a slight upwards curvature, potentially due to a temperature-independent contribution from core diamagnetism (bismuth) or the quartz sample holder. To account for the curvature, we attempted fits utilizing a temperature-independent term, though the result was an nonphysical effective moment of 11 . We opt instead to utilize a limited high-temperature regime (red) for a simpler Curie-Weiss fit (black), which yields a =53 K and a = 9.9. The effective moment agrees well with that expected of Tb3+ (9.72). However, the Weiss temperature is unusually large (53 K), likely affected by the temperature independent contribution, limited temperature range of the fit, and crystal-field effects.
@@
Figure 3(e) presents the zero-field heat capacity (plotted in ) data collected on a single crystal of TbTi3Bi4 down to 100 mK with measurements of LaTi3Bi4 used as a nonmagnetic standard. Two data sets were combined to construct this data set, a conventional helium measurement (200 K – 1.8 K) and a measurement performed with a dilution refrigerator insert (4 K – 100 mK). The data sets did not require substantial adjustment or scaling to overlap in the region from 4 K – 1.8 K after normalization by the sample mass. The two magnetic transitions are evident around 20 K and 2 K. At temperatures below 2 K, a nuclear Schottky peak dominates the heat capacity signal. The peak is still increasing at 0.2 K, which agrees relatively well with a Tb-based Schottky peak ( K).[56] Utilizing LaTi3Bi4 as the nonmagnetic standard (black) and invoking a simple nuclear Schottky model , we can extract the magnetic entropy for TbTi3Bi4.
@@
Figure 3(f) shows the integrated magnetic entropy (green) as a function of temperature. The extracted magnetic contribution to the heat capacity () is also shown (gray) as a schematic reference for the magnetic transition temperatures (not to scale). Interestingly, integrating up to the primary magnetic transition appears to correspond with an entropy release of approximately . Contributions to the magnetic entropy gradually continue up to 100 K, however, consistent with observations in other LnTi3Bi4 materials.[31] By 100 K we recover approximately 90% of the total entropy expected for Tb3+. Considering that the higher temperature heat capacity data tends to be noisier (and that the nonmagnetic subtraction is more prone to error at high temperatures), this is remarkably good agreement with expectations.
We now return to the hysteretic metamagnetic transitions observed in Figure 3(b) and the brief plateau near 2/3 M (Figure 3(b,inset)). Closer inspection of the second metamagnetic transition (H 2.8–3.0 T) revealed a much more complex magnetic landscape than the 1.8 K isothermal curve would suggest. Figure 4(a,d) present our efforts to map the magnetic landscape of TbTi3Bi4, demonstrating a set of isothermal magnetization curves collected in both the increasing (a) and decreasing (d) field directions. As mentioned previously, care must be taken to account for the thermal and magnetic history of TbTi3Bi4 samples. Each isothermal curve was performed by warming the sample above 50 K, and then cooling to the desired temperature under zero-field conditions (ZFC). As such, the magnetic history is erased for each measurement. No ZFC virgin magnetization curves are shown here, but they have been included in the ESI for completeness.
The sharpness of the second transition belies a network of subtle changes in the magnetization. These transitions are quantitatively consistent between samples and between sample batches, suggesting that the phenomenon observed in Figure 3 and Figure 4 are intrinsic to TbTi3Bi4. Taking the field-derivative of the magnetization data shown in (a,d) yields the phase diagrams shown in Figure 4(b,e). Even at this scale, the plethora of phase pockets near 2/3 M is immediately evident, though they are most pronounced at intermediate temperatures from 2 K–15 K.
@@
To better highlight the intricacy of the magnetic landscape near the 2/3 M plateau, we collected higher density data (Figure 4(c,f)) over limited temperature [2 K–18 K] and field [2 T–3.5 T] ranges. The magnetic history was erased between each isothermal curve as described previously. ZFC virgin curves have been omitted, though they have been included in the ESI. While the number of pockets appears to be roughly conserved, the qualitative behavior of the diagram depends strongly on whether the data is collected upon increasing or decreasing field. A series of line cuts at various temperatures (white dashed lines) have been included alongside both phase diagrams to help visualize the phase boundaries. Considering the complexity of the magnetic landscape, neutron diffraction experiments will be invaluable to identify the types and natures of the individual magnetic phases.
Considering the complexity of the magnetic landscape, we were curious whether the features observed in Figure 3 and Figure 4 would have a dramatic influence on the magnetotransport. Cleavable antiferromagnetic materials have recently been a focus in fields like spintronics,[57, 58] and while the low-temperature magnetism of the rare-earth metals may prevent immediate practical application of these materials, there is still insight to be gleaned from the transport data. In particular, probing the connection between complex magnetic phases (via the supporting sublattice) and the transport properties (dominated by the kagome Ti states at E) provides insight into how the sublattices interact.
Figure 5(a) shows the zero-field, in-plane resistivity of TbTi3Bi4 with (8 mA). A simplified schematic of our sample setup is shown in the upper corner of Figure 5(a). While the schematic draws the hexagonal faceted crystal to highlight the [010] facet (edge), the actual sample employed in the measurements was a bar cut from the bulk crystal with a length of 0.4 mm and a cross-sectional area of 0.05 mm2. The crystal quality is good (RRR40) and the zero-field transport is indicative of a relatively high mobility metal. The magnetic transition at 20 K is evident in the resistivity, highlighted by a sharp cusp and subsequent drop. This is perhaps the expected result for ordering from the paramagnetic regime, sharply reducing spin-scattering.
@@
We now apply a magnetic field () to examine the in-plane (transverse) magnetoresistance (MR) of TbTi3Bi4. Figure 5(b) shows the symmetrized transverse MR with and (8 mA) at 1.8 K. As before, we are primarily interested in the magnetotransport under repeated field cycles and have omitted the ZFC virgin curves. Instead, these measurements begin in the fully polarized (9 T) state, ramping to 9 T, and finally back to 9 T. Symmetrized curves corresponding to increasing field direction (in both positive and negative directions) have been indicated by solid purple lines, whereas the decreasing field direction is represented by dashed purple lines. As one would expect from the magnetization results, we observe two sharp transitions in the MR, one for each of the metamagnetic features. The first metamagnetic transition marks a sharp increase in the MR of approximately 30–40%. The second metamagnetic transition marks a sharp decrease in the MR to approximately 50%, which one may expect for a total polarization of the Tb moments. We find the sharp switching between positive and negative MR intriguing in TbTi3Bi4, particularly for a system which is already quite metallic (1Ohm-cm at 1.8 K).
@@
The strong MR response in TbTi3Bi4 at 1.8 K highlights the sensitivity of the transport to the particular magnetic state. Given the complexity of the magnetic phase pockets near the plateau at 2/3 M (see Figure 4(c,f)), we were curious whether the electronic transport would be sensitive to the subtle shifts in the bulk magnetization. Figure 5(c) provides a series of MR curves as a function of temperature. For simplicity and clarity, we are only showing the data under increasing fields, though full field loops are shown in the ESI. The field range corresponding to the 2/3 M plateau (see Figure 3(b)) has been accentuated by the light gray background. Immediately before the transition to negative MR, we observe that the MR cusp becomes a peak – then splits into two/three peaks – merge back together – and then vanish as the system approaches the Néel temperature. These effects occur at the intermediate temperatures consistent with the appearance of the various phase pockets in Figure 4. This further suggests that these are bulk effects corresponding to distinct types of magnetic order.
Utilizing the same sample geometry as Figure 5(a–c), we turned to investigate the Hall effect measurement with and . It is critical to note that there are no metamagnetic transitions with , so both the MR and Hall effect are probing a much simpler magnetic landscape. Furthermore, the effect of hysteresis (and the sensitivity to field and thermal history) vanishes. Figure 5(d) shows the antisymmetrized Hall resistivity for TbTi3Bi4 over two temperature regimes. The top figure shows the Hall response within the antiferromagnetic state (2 K – 18 K), and the bottom figure highlights temperatures within the paramagnetic phase (20 K – 30 K).
As with the MR, full field loops were collected at each temperature, starting at the fully-polarized 7 T state, ramping to 7 T, and finally back to 7 T. No subtractions or corrections have been performed besides antisymmetrization. All data sets shown here exhibit strongly non-linear behavior. A signal that looks curiously like an anomalous Hall signal is evident at 2 K, though the signal persists in shape and magnitude above the magnetic transition temperature. Furthermore, at high temperatures we recover an additional inflection in the data. The signal is also relatively small, leading us to suspect that we may be seeing the effect of multiband transport. Finally, Figure 5(e) presents the transverse MR when and . In the absence of the metamagnetism, no hysteresis in the MR is observed. Further, the system exhibits only strong positive MR (200%) over all fields and temperatures.
Nearly simultaneously with this work, a manuscript focusing on the magnetotransport and ARPES in TbTi3Bi4 were posted.[54] This work complements the results shown here by focusing on the Hall effect when (inducing the metamagnetic transitions). The authors note an record-high anomalous Hall conductivity, nearly an order of magnitude larger than comparable systems.